# Self-Improving Robust Preference Optimization

### v3

@@ 4.2 Optimizing the Robust Generative Policy $\pi$
$\beta\log(Z^{*}(y_{1},x))=\beta(\log(\pi_{\text{ref}}(y_{1}|y_{1},x))-\log(\pi^{*}_{\dagger}(y_{1}|y_{1},x)))-\frac{1}{2}.$
Thus the objective of Eq. (3) can be expressed in terms of $\log(\pi^{*}_{\dagger}(y_{1}|y_{1},x))$ (up to an additive and multiplicative constant) as follows:
@@
$J^{*}(x)\propto\min_{\pi}\mathop{\mathbb{E}}_{y_{1}\sim\pi(.|x)}\left[\log\left(\frac{\pi_{\text{ref}}(y_{1}|y_{1},x)}{\pi^{*}_{\dagger}(y_{1}|y_{1},x)}\right)+D_{\text{KL}}(\pi||\pi_{\text{ref}}|x))\right]\,.$ (9)
@@
where $Z^{*}(x)$ is the normalization factor. Again by taking the logarithm from both side we obtain
$\log(\pi^{*}(y|x))=\log\left(\frac{\pi_{\text{ref}(y|x)}}{\pi_{\text{ref}}(y|y,x)}\pi^{*}_{\dagger}(y|y,x)\right)-\log(Z^{*}(x)).$
Now by collecting terms in Eq. (4) and solving for $\log(\pi^{*}_{\dagger}(y_{2}|y_{1},x)))$ we obtain
$\log(\pi^{*}_{\dagger}(y_{2}|y_{1},x))=\frac{p(y_{2}\succ y_{1}|x)}{\beta}-\log(Z^{*}(y_{1},x)))-\log(\pi_{\text{ref}}(y_{2}|y_{1},x))$ (11)
Now by plugging Eq. (2) into Eq. (10) we deduce
$\pi^{*}(y|x)=\frac{\exp\left(-\log(Z^{*}(y,x))\right)\pi_{\text{ref}}(y|x)}{Z^{*}(x)}.$ (12)
Solving this equation with respect to $\log(Z^{*}(y,x))$ implies
@@
Combining Eq. (11) and Eq. (13) we have for any $y_{1}$ and $y_{2}$
$\frac{p(y_{2}\succ y_{1}|x)}{\beta}-\log\left(\frac{\pi^{*}_{\dagger}(y_{2}|y_{1},x)}{\pi_{\text{ref}}(y_{2}|y_{1},x)}\right)$ $=\log\left(\frac{\pi_{\text{ref}}(y_{1}|x))}{\pi^{*}(y_{1}|x)}\right)-\log(Z^{*}(x)),$
$\frac{p(y_{1}\succ y_{2}|x)}{\beta}-\log\left(\frac{\pi^{*}_{\dagger}(y_{1}|y_{2},x)}{\pi_{\text{ref}}(y_{1}|y_{2},x)}\right)$ $=\log\left(\frac{\pi_{\text{ref}}(y_{2}|x))}{\pi^{*}(y_{2}|x)}\right)-\log(Z^{*}(x)).$
Subtracting these two Equations and collecting terms leads to our key result in which we express the preference $p$ in terms of the optimal self-improvement policy $\pi^{*}_{\dagger}$ and the optimal robust policy $\pi^{*}$.
$p(y_{2}\succ y_{1}|x)=$ $\frac{1}{2}+\frac{\beta}{2}\Bigg{[}\log\left(\frac{\pi^{*}_{\dagger}(y_{2}|y_{1},x)}{\pi_{\text{ref}}(y_{2}|y_{1},x)}\right)-\log\left(\frac{\pi^{*}(y_{1}|x)}{\pi_{\text{ref}}(y_{1}|x))}\right)$ (14)
$-\left(\log\left(\frac{\pi^{*}_{\dagger}(y_{1}|y_{2},x)}{\pi_{\text{ref}}(y_{1}|y_{2},x)}\right)-\log\left(\frac{\pi^{*}(y_{2}|x)}{\pi_{\text{ref}}(y_{2}|x))}\right)\right)\Bigg{]}.$
@@
One may notice the similarity of this result and Equation 6 of DPO paper (Rafailov et al., 2023). Both results express $p(y_{2}\succ y_{1}|x)$ in terms of the optimal policy $\pi^{*}$. However the result of DPO only holds under the assumption that $p$ conforms to the Bradly-Terry model, whereas our result is general and holds for all $p$.
To optimize for $\pi$ and $\pi_{\dagger}$ using (14) we enforce this equation for all $y_{1}$ and $y_{2}$ through following $\ell_{2}$ loss:
$L(\pi,\pi_{\dagger})=$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}y_{1},y_{2}\sim\mu(\cdot|x)\\ x\sim\rho\end{subarray}}\Bigg{[}p(y_{2}\succ y_{1}|x)-\frac{1}{2}-\frac{\beta}{2}\bigg{[}\log\left(\frac{\pi_{\dagger}(y_{2}|y_{1},x)}{\pi_{\text{ref}}(y_{2}|y_{1},x)}\right)-\log\left(\frac{\pi(y_{1}|x)}{\pi_{\text{ref}}(y_{1}|x))}\right)$ (15)
$-\left(\log\left(\frac{\pi_{\dagger}(y_{1}|y_{2},x)}{\pi_{\text{ref}}(y_{1}|y_{2},x)}\right)-\log\left(\frac{\pi(y_{2}|x)}{\pi_{\text{ref}}(y_{2}|x))}\right)\right)\bigg{]}\Bigg{]}^{2}.$
Using the standard properties of $\ell_{2}$-norm to replace $P(y_{2}\succ y_{1}|x)$ with $\mathbf{1}(y_{1}\succ y_{2}|x)$, as $P(y_{2}\succ y_{1}|x)=\mathop{\mathbb{E}}(\mathbf{1}(y_{1}\succ y_{2}|x))$, in the loss of Eq. (15) allows us to derive the following sample loss for SRPO:
@@
$\widehat{L}(\pi,\pi_{\dagger})=$ $\mathop{\mathbb{E}}_{(y_{l},y_{w},x)\sim\mathcal{D}}\Bigg{[}\beta\bigg{[}\log\left(\frac{\pi_{\dagger}(y_{w}|y_{l},x)}{\pi_{\text{ref}}(y_{w}|y_{l},x)}\right)+\log\left(\frac{\pi(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x))}\right)$ (16)
@@
$-\left(\log\left(\frac{\pi_{\dagger}(y_{l}|y_{w},x)}{\pi_{\text{ref}}(y_{l}|y_{w},x)}\right)+\log\left(\frac{\pi(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x))}\right)\right)\bigg{]}-1\Bigg{]}^{2}.$
4.3 Full (Combination) Loss for SRPO
@@
We note that both (16) and (8) are aligned in the sense that both losses are optimizing the same objective of (1). So one can use the convex combination of these two losses as the full loss of SRPO. Also one can use a single LLM (denoted by $\pi$) to represent both $\pi$ and $\pi_{\dagger}$ by exploiting the in-context learning power of LLMs (Brown et al., 2020) such that $\pi_{\dagger}(y^{\prime}|y,x)=\pi(y^{\prime}|y,x)$. So for every $\alpha\in[0,1]$ we define the full sample loss of SRPO as follows
@@
5:     Estimate $\nabla_{\theta}\widehat{L}_{\alpha}(\pi_{\theta})$ from Eq. (17) using minibatch $B$ as the dataset
@@ 5 Robustness of SRPO
In the case of both IPO and DPO the analytical solution is already well-established and analyzed for both algorithms (Azar et al., 2023; Rafailov et al., 2023; Tang et al., 2024). In particular the optimal solution for both IPO and DPO can be expressed explicitly in terms of the soft-max of the expected preference as follows (Azar et al., 2023):
@@
$\pi^{*}(y|x)=\frac{\exp\left(\frac{\mathop{\mathbb{E}}_{y^{\prime}\sim\mu(\cdot|x)}(\Psi(p(y\succ y^{\prime}|x)))}{\beta}\right)\pi_{\text{ref}}(y|x)}{Z^{*}(x)},$ (18)
@@
with the choice of $\Psi=I(\cdot)$ and $\Psi=\sigma^{-1}(\cdot)$ for IPO and DPO, respectively, where $\sigma^{-1}$ denotes the inverse-sigmoid (logit function). Thus, based on (18), we can see that the solution for both IPO and DPO has strong dependency on $\mu$ in the form of expected preference under the distribution $\mu$. Thus it may not be robust to changes in $\mu$. This dependency on $\mu$ can be especially problematic when we evaluate the model on out-of-distribution tasks where the desired behavior is very different from $\mu$ and the expected preference under the distribution $\mu$ is not a good measure of performance. SRPO solution on the other hand has no dependency on the behavior policy $\mu$: from (2) we observe that the optimal self-improvement policy $\pi^{*}_{\dagger}$ is independent of $\mu$ and, unlike DPO and IPO cases, is expressed in terms of softmax of $P(y_{2}\succ y_{1}|x)$ for any pair of completions $(y_{1},y_{2})$. Also the 00-revision policy $\pi^{*}$ is also completely independent of $\mu$ as it is evident from (10) (i.e., it is proportional to $\pi^{*}(y|y,x)$ which itself is independent of $\mu$). Thus, from a mathematical point of view, SRPO provides a robust solution for the problem of direct preference optimization that does not depend on the behavior policy $\mu$.
To illustrate further the differences between SRPO and DPO/IPO with regard to robustness to $\mu$ we consider a simple bandit example. For simplicity we assume there is no context $x$, i.e., we are in a standard bandit setting. Consider the simple case where we have 3 actions (completions) $y_{0}$, $y_{1}$ and $y_{2}$, for which the preference model is given as follows:
$P=\begin{pmatrix}0.5&0.99&0.3\\ 0.01&0.5&0.25\\ 0.7&0.75&0.5\end{pmatrix},$
in which $p(y_{i}\succ y_{j})=P_{ij}$. In this case $y_{2}$ is clearly the preferred outcome as it dominates both $y_{1}$ and $y_{0}$ by probability larger than $0.5$. On the other hand, if we only consider preference with respect to $y_{1}$ then arm $y_{0}$ is a better outcome as it is preferred to $y_{1}$ with higher probability than $y_{2}$. Therefore, if a preference optimization method is robust to changes in $\mu$ one might expect that the optimal solution should be independent of the frequency of $y_{2}$ in the preference dataset (i.e., $\mu(y_{2})$).
@@
To test this hypothesis we consider two synthetic dataset of actions generated from distributions $\mu_{0}$ and $\mu_{1}$: We set $\mu_{0}$ to be a uniform behavior policy ($\mu_{0}(y_{0})=\mu_{0}(y_{1})=\mu_{0}(y_{2})=\frac{1}{3}$) and $\mu_{1}$ skewed towards $y_{1}$ ($\mu_{1}(y_{1})=0.7,\mu_{1}(y_{0})=\mu_{1}(y_{2})=0.15$). We then generate a dataset of $10000$ pairs from $\mu_{0}$ and $\mu_{1}$ and rate them according to the preference model $p$ (for any pair $(y_{1},y_{2})$ we assign the preference by sampling from $p(y_{1}\succ y_{2})$, that is $y_{1}$ is preferred to $y_{2}$ with probability $p(y_{1}\succ y_{2})$). This provides us with two dataset of rated completions $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$ for $\mu_{0}$ and $\mu_{1}$. We then use these two datasets to train the policy $\pi$ using SRPO, DPO and IPO using a simple Adam optimizer. In the case of IPO and DPO we optimize only the 00-revision policy $\pi(y)$ where as for SRPO we also optimize the self-improvement policy $\pi(y|y^{\prime})$ as well. We set the regularization constant $\beta$ for all methods to $1$. We consider a uniform distribution $\pi_{\text{ref}}(y)=1/3$ for all algorithms and all $y$s. In the case of SRPO we set the self-improvement reference policy $\pi_{\text{ref}}(y|y^{\prime})=1/3$ for all $y$ and $y^{\prime}$. Also for SRPO we set the combination coefficient $\alpha=0$ for simplicity.
@@
We observe that in the case of using uniform $\mu_{0}$ as a behavior policy all methods do the right thing and their policies converge to solutions in which $y_{2}$ dominates $y_{1}$ and $y_{0}$ (Fig. 1(a)). However, when we use the behavior policy $\mu_{1}$ which is skewed towards $y_{1}$, both DPO and IPO converge to a solution in which $y_{0}$ dominates $y_{1}$ and $y_{2}$, while the policy of SRPO remains intact (Fig. 1(b)). Notice that the SRPO policy is slightly different in both cases. This is to be expected, we are in a finite data setting, and the sampling distribution will have some influence on the empirical preference model (defining the empirical solution of SRPO).
(a) SRPO vs IPO and DPO for uniform behavior $\mu_{0}$.
(b) SRPO vs. IPO and DPO for skewed behavior $\mu_{1}$.
@@
Figure 1: Learned action probabilities for the synthetic example. SRPO always chooses the correct arm regardless of skew in the behaviour policy $\mu$, while both IPO and DPO are effected by the skew as portrayed in Fig. (1(b)).
6 Related works
Our work lies in offline preference optimization, a vivid area of research since the introduction of DPO (Rafailov et al., 2023). Some of the core concepts of this research topic was generalized and formalized by Azar et al. (2023). In particular they characterized the underlying optimal solution for a generic preference optimization objective and introduced IPO for addressing some of the related shortcomings of DPO. SLiC-HF (Zhao et al., 2023a) was introduced around the same time, from a less RL-centric point of view. All these approaches have been abstracted later by Tang et al. (2024), the general recipe being to build a contrastive loss function from a convex classification function and to make use of the analytical solution of the RL problem to learn directly the policy. A common underlying assumption is that the related RL problem is KL-regularized. This has been generalized to more general $f$-divergences by Wang et al. (2023). These are just a few among many works on direct alignment. However, they all share the fact of not considering self-improvement policies, contrary to SRPO. This has a strong incidence on the related solution concept, making SRPO the sole direct alignment approach being robust to the sampling distribution $\mu$, as showcased in Sec. 5.
Some similarities also exist between SRPO and Nash-MD (Munos et al., 2023). Indeed, if in Eq. 1 we replace the self-improvement policy $\pi_{\dagger}(\cdot|y,x)$ by a classic policy $\pi(\cdot|x)$, then we obtain the saddle-point optimization problem that Nash-MD solves. However, considering a self-improvement policy is a core contribution of our work, and it is not anecdotal. From a technical viewpoint, this is critical for simplifying the minimax problem of Eq. (1) and obtaining a simple offline optimization problem. NashMD on the other hand adapts algorithms from the game-theory literature and can only be solved online with all the stability issues of online methods and large inference costs. From practical point of view self-improvement provides a boost in performance by refining the original generations of LLM. The feature that Nash-MD is missing. Finally, even though the Nash equilibrium of Nash-MD does not depend on the sampling distribution $\mu$, it relies on a learned reward function, with the possible associated caveats mentioned earlier, which is not the case of SRPO.
Our work is also obviously related to the concept of chain of thoughts (Wei et al., 2023; Yao et al., 2024), self-improvement (Huang et al., 2022) and self-refining LLMs (Madaan et al., 2024). However, it is very often used as a way of prompting a model to obtain better results, and less often as a component of a learning paradigm (Liu et al., 2023; Huang et al., 2022). To our best knowledge, we propose the first approach that combines training self-improvement LLMs and offline preference optimization through a single supervised objective, moreover in a theoretically grounded manner and showing the robustness to $\mu$.

### v4

@@ 4.2 Proof of Main Result
$\pi^{*}_{\dagger}(y^{\prime}|y,x)=\frac{\exp\left(\frac{p(y^{\prime}\succ y|x)}{\beta}\right)\pi_{\text{ref}}(y^{\prime}|y,x)}{Z^{*}(y,x)},$ (5)
where $Z^{*}(y,x)$ is the normalization factor. One can easily show that by plugging $\pi^{*}_{\dagger}$ in the objective function of (1) we obtain:
$J^{*}(x)=\min_{\pi}\mathop{\mathbb{E}}_{y\sim\pi(.|x)}\left[\beta(\log(Z^{*}(y,x))+D_{\text{KL}}(\pi||\pi_{\text{ref}}|x))\right].$ (6)
Now by solving (5) with respect to $p(y^{\prime}\succ y|x)$ we obtain
$p(y^{\prime}\succ y|x)=\beta(\log(\pi^{*}_{\dagger}(y^{\prime}|y,x))-\log(\pi_{\text{ref}}(y^{\prime}|y,x))+\beta\log(Z^{*}(y,x))).$ (7)
4.2.1 Optimizing the Self-Improvement Policy $\pi_{\dagger}$
We now prove that minimizing the objective of (3) gives us the optimal self-improvement policy (solution to maximization in the objective of (1)). We first notice that using the convention $p(y\succ y|x)=\frac{1}{2}$ (7) implies
$\frac{1}{2}=\beta(\log(\pi^{*}_{\dagger}(y|y,x))-\log(\pi_{\text{ref}}(y|y,x)))+\beta\log(Z^{*}(y,x)).$ (8)
Now by subtracting (7) from (8) we derive
@@ 4.2.1 Optimizing the Self-Improvement Policy $\pi_{\dagger}$
$p(y^{\prime}\succ y|x)=\frac{1}{2}+\beta\left[\log\left(\frac{\pi^{*}_{\dagger}(y^{\prime}|y,x)}{\pi_{\text{ref}}(y^{\prime}|y,x)}\right)-\log\left(\frac{\pi^{*}_{\dagger}(y|y,x)}{\pi_{\text{ref}}(y|y,x)}\right)\right].$ (9)
This is our first key result that expresses preference $p(y^{\prime}\succ y|x)$ in terms of the optimal self-improvement policy $\pi^{*}_{\dagger}$. So we enforce this equation for all $y$ and $y^{\prime}$ through minimizing $\ell_{2}$ loss of (3) which concludes the proof for the self-improvement policy.
4.2.2 Joint Optimization of the Robust Generative Policy $\pi$ and the improvement policy $\pi_{\dagger}$
In this section, we show that to optimize the min-max objective of (1) (i.e., solving for $\pi$ and $\pi_{\dagger}$) one can optimize the loss of (4). We start by collecting terms in (8) which implies $\beta\log(Z^{*}(y,x))=\beta(\log(\pi_{\text{ref}}(y|y,x))-\log(\pi^{*}_{\dagger}(y|y,x)))-\frac{1}{2}$. Thus, the objective of (6) can be expressed in terms of $\log(\pi^{*}_{\dagger}(y|y,x))$ (up to an additive and multiplicative constant) as follows:
$J^{*}(x)\propto\min_{\pi}\mathop{\mathbb{E}}_{y\sim\pi(.|x)}\left[\log\left(\frac{\pi_{\text{ref}}(y|y,x)}{\pi^{*}_{\dagger}(y|y,x)}\right)+D_{\text{KL}}(\pi||\pi_{\text{ref}}|x))\right]\,.$ (10)
@@ 4.2.2 Joint Optimization of the Robust Generative Policy $\pi$ and the improvement policy $\pi_{\dagger}$
where $Z^{*}(x)$ is the normalization factor. Again by taking the logarithm from both side we obtain $\log(\pi^{*}(y|x))=\log\left(\frac{\pi_{\text{ref}}(y|x)}{\pi_{\text{ref}}(y|y,x)}\pi^{*}_{\dagger}(y|y,x)\right)-\log(Z^{*}(x))$. Now by collecting terms in (7) and solving for $\log(\pi^{*}_{\dagger}(y^{\prime}|y,x))$ we obtain
$\log(\pi^{*}_{\dagger}(y^{\prime}|y,x))=\frac{p(y^{\prime}\succ y|x)}{\beta}-\log(Z^{*}(y,x))-\log(\pi_{\text{ref}}(y^{\prime}|y,x))$ (12)
Now by plugging (5) into (11) we deduce $\pi^{*}(y|x)=\frac{\exp\left(-\log(Z^{*}(y,x))\right)\pi_{\text{ref}}(y|x)}{Z^{*}(x)}.$ Solving this equation with respect to $\log(Z^{*}(y,x))$ implies
@@
Combining (12) and (13) for any $y$ and $y^{\prime}$ we have $\frac{p(y^{\prime}\succ y|x)}{\beta}-\log\left(\frac{\pi^{*}_{\dagger}(y^{\prime}|y,x)}{\pi_{\text{ref}}(y^{\prime}|y,x)}\right)=\log\left(\frac{\pi_{\text{ref}}(y|x)}{\pi^{*}(y|x)}\right)-\log(Z^{*}(x)),$ and also $\frac{p(y\succ y^{\prime}|x)}{\beta}-\log\left(\frac{\pi^{*}_{\dagger}(y|y^{\prime},x)}{\pi_{\text{ref}}(y|y^{\prime},x)}\right)=\log\left(\frac{\pi_{\text{ref}}(y^{\prime}|x)}{\pi^{*}(y^{\prime}|x)}\right)-\log(Z^{*}(x))$. Subtracting these two equations and collecting terms leads to our key result in which we express the preference $p$ in terms of the self-improvement policy $\pi^{*}_{\dagger}$ and the robust policy $\pi^{*}$.
$p(y^{\prime}\succ y|x)=$ $\frac{1}{2}+\frac{\beta}{2}\Bigg{[}\log\left(\frac{\pi^{*}_{\dagger}(y^{\prime}|y,x)}{\pi_{\text{ref}}(y^{\prime}|y,x)}\right)+\log\left(\frac{\pi^{*}(y|x)}{\pi_{\text{ref}}(y|x)}\right)$ (14)
$-\left(\log\left(\frac{\pi^{*}_{\dagger}(y|y^{\prime},x)}{\pi_{\text{ref}}(y|y^{\prime},x)}\right)+\log\left(\frac{\pi^{*}(y^{\prime}|x)}{\pi_{\text{ref}}(y^{\prime}|x)}\right)\right)\Bigg{]}.$
@@
One may notice the similarity of this result and Eq. 6 of DPO paper (Rafailov et al., 2023). Both results express $p(y^{\prime}\succ y|x)$ in terms of the optimal policy $\pi^{*}$. However, the result of DPO only holds under the assumption that $p$ conforms to the Bradly-Terry model, whereas our result is general and holds for all $p$.
To optimize for $\pi$ and $\pi_{\dagger}$ using (14) we enforce this equation for all $y$ and $y^{\prime}$ through minimizing the expected $\ell_{2}$ loss of (4) which concludes the proof on joint optimization of the generative policy $\pi$ and the self-improvement policy through (4).
4.2.3 Putting all Together in a Single Combination Loss
We observe that the losses defined in (4) and (3) are inherently aligned, as both aim to optimize the same overall objective given in (1). This alignment allows us to construct a combined loss function for SRPO by using a convex combination of these two losses. Specifically, for any $\alpha\in[0,1]$, minimizing the convex combination of the losses in (4) and (3) is equivalent to directly optimizing the objective in (1). Mathematically, the full loss of SRPO can be expressed as:
$\mathcal{L}_{\text{{SRPO}}}=\alpha\cdot\mathcal{L}_{\text{L2}}+(1-\alpha)\cdot\mathcal{L}_{\text{self-improve}},$
where $\mathcal{L}_{\text{L2}}$ corresponds to the loss in (4) and $\mathcal{L}_{\text{self-improve}}$ corresponds to the loss in (3). By minimizing this combined loss, we ensure that the optimization process aligns with the original objective in (1) regardless of the value of $\alpha$.
This formulation not only unifies the two loss functions within a single framework but also provides flexibility in adjusting their relative contributions during training. This completes the proof.
4.3 Deriving the Sample Loss for SRPO
Calculating/optimizing the expected loss of (17) is often not practical as we often have no direct access to the behavior policy $\mu$. Moreover, the space of completions $y$ is often very large (even infinite). So calculating the full expectation is often impractical. Instead one can estimate the expected loss $L_{\alpha}$ of (17) using a preference dataset $\mathcal{D}$ with a sample loss $\widehat{L}_{\alpha}$ and optimize this loss instead. We now focus on deriving the sample loss $\widehat{L}_{\alpha}$ for SRPO. Using the standard properties of $\ell_{2}$-norm to replace $p(y^{\prime}\succ y|x)$ with $\mathbf{1}(y^{\prime}\succ y|x)$, as $p(y^{\prime}\succ y|x)=\mathop{\mathbb{E}}[\mathbf{1}(y^{\prime}\succ y|x)]$, in the objective of (4) allows us to derive the following sample loss for the improvement model:
@@ 4.3 Deriving the Sample Loss for SRPO
$\widehat{L}_{\dagger}(\pi_{\dagger})=$ $\mathop{\mathbb{E}}_{\begin{subarray}{c}(y_{l},y_{w},x)\sim\mathcal{D}\end{subarray}}\left[\frac{1}{2}-\beta\left[\log\left(\frac{\pi_{\dagger}(y_{w}|y_{l},x)}{\pi_{\text{ref}}(y_{w}|y_{l},x)}\right)-\log\left(\frac{\pi_{\dagger}(y_{l}|y_{l},x)}{\pi_{\text{ref}}(y_{l}|y_{l},x)}\right)\right]\right]^{2}$ (15)
@@
$+\mathop{\mathbb{E}}_{\begin{subarray}{c}(y_{l},y_{w},x)\sim\mathcal{D}\end{subarray}}\left[\frac{1}{2}-\beta\left[\log\left(\frac{\pi_{\dagger}(y_{w}|y_{w},x)}{\pi_{\text{ref}}(y_{w}|y_{w},x)}\right)-\log\left(\frac{\pi_{\dagger}(y_{l}|y_{w},x)}{\pi_{\text{ref}}(y_{l}|y_{w},x)}\right)\right]\right]^{2}.$
Using the standard properties of $\ell_{2}$-norm to replace $p(y^{\prime}\succ y|x)$ with $\mathbf{1}(y^{\prime}\succ y|x)$, as $p(y^{\prime}\succ y|x)=\mathop{\mathbb{E}}[\mathbf{1}(y^{\prime}\succ y|x)]$, in the loss of (4) allows us to derive the following sample loss for both the generative policy and the improvement model:
$\widehat{L}(\pi,\pi_{\dagger})=$ $\mathop{\mathbb{E}}_{(y_{l},y_{w},x)\sim\mathcal{D}}\Bigg{[}\beta\bigg{[}\log\left(\frac{\pi_{\dagger}(y_{w}|y_{l},x)}{\pi_{\text{ref}}(y_{w}|y_{l},x)}\right)+\log\left(\frac{\pi(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}\right)$ (16)
$-\left(\log\left(\frac{\pi_{\dagger}(y_{l}|y_{w},x)}{\pi_{\text{ref}}(y_{l}|y_{w},x)}\right)+\log\left(\frac{\pi(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right)\bigg{]}-1\Bigg{]}^{2}.$
@@
looseness=-1 Furthermore, one can use a single LLM (denoted by $\pi$) to represent both $\pi$ and $\pi_{\dagger}$ by exploiting the in-context learning power of LLMs (Brown et al., 2020) such that $\pi_{\dagger}(y^{\prime}|y,x)=\pi(y^{\prime}|y,x)$. So for we define the full sample loss of SRPO as the convex combination of (16) and (15):
@@
5:     Estimate $\nabla_{\theta}\widehat{L}_{\alpha}(\pi_{\theta})$ from (17) using minibatch $B$ as the dataset
@@ 5 Robustness of SRPO
In the case of both IPO and DPO the analytical solution is already well-established and analyzed for both algorithms (Azar et al., 2023; Rafailov et al., 2023; Tang et al., 2024). In particular, the optimal solution for both IPO and DPO can be expressed explicitly in terms of the soft-max of the expected preference as follows (Azar et al., 2023):
@@
$\pi^{*}(y|x)=\frac{\exp\left(\beta^{-1}\mathop{\mathbb{E}}_{y^{\prime}\sim\mu(\cdot|x)}[\Psi(p(y\succ y^{\prime}|x))]\right)\pi_{\text{ref}}(y|x)}{Z^{*}(x)},$ (18)
@@
with the choice of $\Psi=I(\cdot)$ and $\Psi=\sigma^{-1}(\cdot)$ for IPO and DPO, respectively, where $\sigma^{-1}$ denotes the inverse-sigmoid (logit function). Thus, based on (18), we can see that the solution for both IPO and DPO has strong dependency on $\mu$ in the form of expected preference under the distribution $\mu$. Thus it may not be robust to changes in $\mu$. This dependency on $\mu$ can be especially problematic when we evaluate the model on out-of-distribution tasks where the desired behavior is very different from $\mu$ and the expected preference under the distribution $\mu$ is not a good measure of performance. SRPO solution on the other hand has no dependency on the behavior policy $\mu$: from (5) we observe that the optimal self-improvement policy $\pi^{*}_{\dagger}$ is independent of $\mu$ and, unlike DPO and IPO cases, is expressed in terms of softmax of $p(y^{\prime}\succ y|x)$ for any pair of completions $(y,y^{\prime})$. Also the 00-revision policy $\pi^{*}$ is also completely independent of $\mu$ as it is evident from (11) (i.e., it is proportional to $\pi^{*}(y|y,x)$ which itself is independent of $\mu$). Thus, from a mathematical point of view, SRPO provides a robust solution for the problem of direct preference optimization that does not depend on the behavior policy $\mu$. To illustrate the differences between SRPO and DPO/IPO regarding robustness to $\mu$, we consider a simple bandit example. For simplicity, we assume there is no context $x$. Consider the simple case where we have 3 completions $y_{0}$, $y_{1}$, and $y_{2}$, for which the preference model is $P=\begin{pmatrix}0.5&0.99&0.3\\ 0.01&0.5&0.25\\ 0.7&0.75&0.5\end{pmatrix}.$ To test this hypothesis we consider two synthetic dataset of actions generated from distributions $\mu_{0}$ and $\mu_{1}$: We set $\mu_{0}$ to be a uniform behavior policy
@@
($\mu_{0}(y_{0})=\mu_{0}(y_{1})=\mu_{0}(y_{2})=1/3$) and $\mu_{1}$ skewed towards $y_{1}$ ($\mu_{1}(y_{1})=0.7,\mu_{1}(y_{0})=\mu_{1}(y_{2})=0.15$). We then generate a dataset of $10,000$ pairs from $\mu_{0}$ and $\mu_{1}$ and rate them according to the preference model $p$ (for any pair $(y,y^{\prime})$ we assign the preference by sampling from $p(y\succ y^{\prime})$, that is $y$ is preferred to $y^{\prime}$ with probability $p(y\succ y^{\prime})$). This provides us with two dataset of rated completions $\mathcal{D}_{0}$ and $\mathcal{D}_{1}$ for $\mu_{0}$ and $\mu_{1}$. We then use these two datasets to train the policy $\pi$ using SRPO, DPO and IPO using a simple Adam optimizer. In the case of IPO and DPO we optimize only the 00-revision policy $\pi(y)$ where as for SRPO we also optimize the self-improvement policy $\pi(y|y^{\prime})$ as well. We set the regularization constant $\beta$ for all methods to $1$. We consider a uniform distribution $\pi_{\text{ref}}(y)=1/3$ for all algorithms and all $y$s. In the case of SRPO we set the self-improvement reference policy $\pi_{\text{ref}}(y|y^{\prime})=1/3$ for all $y$ and $y^{\prime}$. Also for SRPO we set the combination coefficient $\alpha=0$ for simplicity.
@@
We observe that in the case of using uniform $\mu_{0}$ as a behavior policy, all methods perform as expected, converging to solutions where $y_{2}$ dominates $y_{1}$ and $y_{0}$ (Fig. 2(a)). However, when the behavior policy $\mu_{1}$ is skewed toward $y_{1}$, both DPO and IPO converge to a solution where $y_{0}$ dominates $y_{1}$ and $y_{2}$, while the policy of SRPO remains unchanged (Fig.2(b)). Notably, the SRPO policy differs slightly between the two cases, which is expected given the finite data setting, where the sampling distribution influences the empirical preference model." However, when we use the behavior policy $\mu_{1}$, which is skewed towards $y_{1}$, both DPO and IPO converge to a solution where $y_{0}$ dominates $y_{1}$ and $y_{2}$, while the policy of SRPO remains unchanged (Fig.2(b)). Notably, the SRPO policy differs slightly between the two cases, which is expected given the finite data setting, where the sampling distribution influences the empirical preference model."
to a solution in which $y_{0}$ dominates $y_{1}$ and $y_{2}$, while the policy of SRPO remains intact (Fig. 2(b)). Notice that the SRPO policy is slightly different in both cases. This is to be expected, we are in a finite data setting, and the sampling distribution will have some influence on the empirical preference model.
(a) SRPO vs IPO and DPO for uniform $\mu_{0}$.
(b) SRPO vs. IPO and DPO for skewed $\mu_{1}$.
@@
Figure 2: Learned action probabilities for the synthetic example. SRPO always chooses the correct arm regardless of skew in $\mu$, while both IPO and DPO are effected by the skew (Fig. (2(b))).